# Anytime-valid, Bayes-assisted, Prediction-Powered Inference

### v1

@@ 6.2 Real data
We evaluate our method on three mean estimation tasks, using the flights, forest, and galaxies datasets, which are described in the supplementary material. Each dataset provides label/prediction pairs $\{(Y_{i},f(X_{i}))\}_{i=1}^{N+n}$, which we randomly split into a set of size $n$, which is used to simulate a labelled data stream, and an unlabelled set of size $N$. Figure 3 compares classical and PPI++ AsympCS procedures on these datasets in terms of average interval volume and cumulative miscoverage rate as $n$ varies. PPI methods consistently yield smaller regions, while maintaining reliable coverage. Moreover, Bayes-assisted approaches further improve efficiency for moderate labelled sample sizes.
Figure 3: Real data study. The top and bottom rows show the average interval volume and cumulative miscoverage rate over $1000$ repetitions for the flights, forest, and galaxies datasets.
@@ 7 Discussion
The proposed approach, like standard PPI, involves certain practical considerations. In general settings beyond the mean and linear regression cases, it typically requires constructing a grid of values for $\theta$. When $\eta_{t}$ is not available in closed form – as with a Student-$t$ prior – numerical integration may be needed. The current development focuses on scalar parameter values $\theta$, though extensions of AsympCS to multivariate settings are available (Waudby-Smith et al., 2024b, §B.10), and it would be interesting to explore whether Bayes-assisted variants can be similarly generalised. The confidence sequences derived in this work are asymptotically valid under mild, nonparametric assumptions. A promising direction for future work is the development of nonasymptotic, nonparametric Bayes-assisted confidence sequences under stricter assumptions, such as bounded means, building on the work of Waudby-Smith and Ramdas (2024).
@@
The supplementary material is organised as follows. Appendix S1 gives additional background on strong laws and couplings, nonnegative martingales and the method of mixtures. Appendix S2 states secondary results and their proofs. Appendix S3 presents the proofs of the main theorems and propositions. Appendix S4 derives tighter convergence rates for the AsympCS introduced in the main text, under stronger moment conditions. Appendix S5 extends the results to the multivariate setting. Appendix S6 gives specific expressions for the case of mean estimation. Appendix S7 details the experimental setup used in the main text. Appendix S8 presents further experiments. Finally, Appendix S9 introduces a parameter-free, non-assisted, AsympCS , and provides some additional comparisons.
For clarity, all sections, theorems, propositions, and lemmas in the supplementary material are prefixed with “S” to distinguish them from those in the main text.
@@
S1.1 Iterated logarithm and Marcinkiewicz-Zygmund strong laws
@@
(Iterated Logarithm Law (Durrett, 2019, Theorem 8.5.2)) Let $(Y_{t})_{t\geq 1}$ be i.i.d. random variables with zero mean and unit variance. Let $S_{t}=\sum_{i=1}^{t}Y_{i}$. Then,
@@
which implies
$\left|\frac{S_{t}}{t}\right|=O\left(\sqrt{\frac{\log\log t}{t}}\right)\text{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ a.s. as }t\to\infty.$
Theorem S2.
(Marcinkiewicz-Zygmund strong law of large numbers (Durrett, 2019, Theorem 2.5.12)) Let $(Y_{t})_{t\geq 1}$ be i.i.d. random variables with zero mean and $\mathbb{E}|Y_{1}|^{p}<\infty$ for some $21<p<$ 2. Let $S_{t}=\sum_{i=1}^{t}X_{i}$. Then,
$\frac{S_{t}}{t^{1/p}}\rightarrow 0\text{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ a.s. as }t\to\infty.$
S1.2 Strong approximations
The following strong invariance result, attributed to Strassen (1964) and stated in Kallenberg (2002, Theorem 14.6), shows that the partial sums of i.i.d. random variables can be approximated almost surely by a Brownian motion path.
Theorem S3 (Strassen’s strong approximation Strassen (1964)).
Let $(Y_{t})_{t\geq 1}$ be i.i.d. random variables with mean 0 and variance 1, and write $S_{t}=\sum_{i=1}^{t}Y_{i}$. Then there exists a standard Brownian motion $B$ such that
$S_{t}-B_{t}=o(\sqrt{t\log\log t})\text{ a.s. as }t\to\infty.$
Strassen’s coupling holds under finite variance assumption. The bound can be tightened under additional moment assumptions, as shown by Kolmos, Major and Tusnady (KMT) Komlós et al. (1975); Major (1976). The following theorem is from Csörgö and Hall (1984, Theorem 3.2).
Theorem S4 (KMT strong coupling Komlós et al. (1975); Major (1976)).
$S_{t}-B_{t}=o(t^{1/q})\text{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ a.s.\leavevmode\nobreak\ as\leavevmode\nobreak\ }t\rightarrow\infty.$
KMT strong coupling has been extended by Einmahl (1987) to random vectors (see also (Waudby-Smith et al., 2024a, Section B11)).
Theorem S5.
(Multivariate KMT strong coupling Einmahl (1987)) Let $(Y_{t})_{t\geq 1}$ be i.i.d. random vectors in $\mathbb{R}^{d}$ with zero mean, covariance matrix $\Sigma$, and such that $\mathbb{E}||Y_{1}||^{q}<\infty$ for some $q>2$. Let $S_{t}=\sum_{i=1}^{t}Y_{i}$. Then, there exists a standard multivariate Brownian motion B such that,
$\Sigma^{-{1/2}}S_{t}-B_{t}=o(t^{1/q})\text{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ a.s.\leavevmode\nobreak\ as\leavevmode\nobreak\ }t\rightarrow\infty.$
S1.3 Nonnegative supermartingale and Ville’s inequality
@@
$M=(M_{t})_{t\geq 1}$ is a nonnegative supermartingale (NSM) with respect to the adapted filtration $(\mathcal{F}_{t})_{t\geq 1}$ if $M_{t}\geq 0$ a.s, $\mathbb{E}[M_{t}]<\infty$ for all $t\geq 1$, and
[…]
@@
Proposition S7 (Ville’s inequality (Ville, 1939)).
@@
Proposition S9 (Lemma B1, Chugg et al. (2023)).
Let $\{(M_{t}(\mu))_{t\in\mathbb{N}},\mu\in\Theta\}$ be a family of (super)martingales on a filtered probability space $(\Omega,\mathcal{A},(\mathcal{F}_{t})_{t\in\mathbb{N}},\Pr)$, indexed by $\mu$ in a measurable space $(\Theta,\mathcal{B})$, such that
@@
    each $M_{t}(\mu)$ is $\mathcal{F}_{t}\otimes\mathcal{B}$-measurable; and
@@
    each $\mathbb{E}\left[M_{t}(\mu)\mid\mathcal{F}_{t-1}\right]$ is $\mathcal{F}_{t-1}\otimes\mathcal{B}$-measurable.
Let $\pi$ be a finite measure on $(\Theta,\mathcal{B})$ such that for all $n$,
$\Pr\otimes\pi\text{-almost everywhere }M_{t}(\mu)\geq 0,\quad\text{ or }\quad\mathbb{E}_{\mu\sim\pi}\mathbb{E}\left[\left|M_{t}(\mu)\right|\right]<\infty$
Then the mixture $\left(\widetilde{M}_{t}\right)_{t\in\mathbb{Z}}$, where $\widetilde{M}_{t}=\mathbb{E}_{\mu\sim\pi}M_{t}(\mu)$, is also a (super)martingale.
[…]
The method of mixtures dates back at least to Ville (1939) and was developed in the context of sequential analysis by Wald (1945). It was then systematised and popularised by Darling and Robbins in the late 1960s, by Robbins and Siegmund in a series of papers culminating in Robbins (1970), and by Lai (1976). The method of mixtures has found many applications, including confidence sequences (Lai, 1976; Balsubramani and Ramdas, 2015; Kaufmann and Koolen, 2021; Howard et al., 2021; Waudby-Smith et al., 2024b), PAC-Bayes analysis (Haddouche and Guedj, 2022; Chugg et al., 2023), anytime-valid testing (Ramdas et al., 2023), and A/B testing (Johari et al., 2017), to name but a few.
@@
The following proposition follows from Strassen’s strong approximation (see Theorem S3). It will be used in the proofs of Theorem 3.5 and Proposition 4.2.

### v2

@@ 6.2 Real data
We evaluate our method on several real-world datasets, which are described in Section˜S6.2. While each dataset is, in principle, static (providing label/prediction pairs $\{(Y_{i},f(X_{i}))\}_{i=1}^{N+n_{1}}$), we simulate an online setting akin to Section˜6.1 by randomly splitting the data into a labelled set of size $n_{1}$, which serves as a labelled data stream, and an unlabelled set of size $N$.
Figure˜3 compares classical and PPI++ AsympCS procedures on the flights, forest, and galaxies datasets, where the goal is mean estimation.
Figure 3: Mean estimation. The top and bottom rows show the average interval volume and cumulative miscoverage rate over $1000$ repetitions for the flights, forest, and galaxies datasets.
By taking advantage of the unlabelled data, PPI methods consistently yield smaller regions than the classical counterpart, while maintaining reliable coverage. Moreover, Bayes-assisted approaches further improve efficiency for moderate labelled sample sizes, as the quality of the predictions is generally high in these datasets.
Figure˜S11 reports results for three additional estimation tasks: linear regression (census), logistic regression (healthcare), and quantile estimation (genes). For the first two tasks, the same conclusions as for mean estimation hold: PPI methods consistently outperform classical inference, with Bayes-assisted approaches providing an additional efficiency boost. For the quantile estimation task, non-assisted PPI still improves over classical inference by leveraging the machine learning predictions; however, the Bayes-assisted methods yield larger regions than the other approaches, reflecting lower prediction quality in this dataset.
@@ 7 Discussion
We extended the PPI framework to the sequential setting via asymptotic confidence sequences, which further allow for seamless integration of prior information about the quality of the auxiliary predictions. However, several directions merit further investigation. The results developed here are for scalar parameter values $\theta$. Extensions to multivariate settings are discussed in Appendix˜S4, building on earlier work by Waudby-Smith et al. [15, §B.10]. In the non-assisted case, we focused on asymptotic confidence sequences of the form (6), but other options are possible. In particular, as discussed in Appendix˜S8, the parameter-free CS proposed by Wang and Ramdas [19], which is based on an improper prior, may be used as an exact reference CS in place of Equation˜6.
The AsympCS derived in this paper are asymptotically valid for i.i.d. data under mild, nonparametric assumptions. Promising directions include extensions to non-i.i.d. observations, as well as the development of nonasymptotic, nonparametric Bayes-assisted confidence sequences under stricter assumptions (e.g., bounded means), building on the work of Waudby-Smith and Ramdas [15]. In the non-assisted case, the parameter $\rho$ was assumed to be fixed. Waudby-Smith et al. [15, §2.5] considered delayed-start sequences $\mathcal{C}_{\alpha,t}(m)$ that may depend on the start time $m$; this includes allowing the tuning parameter $\rho$ to depend on $m$. Their asymptotic Type-I error control result, derived under assumptions similar to those used here, also applies in our setting. Another interesting direction would be to adapt similar ideas to the Bayes-assisted construction.
PPI AsympCS procedures share the computational considerations of their fixed-time counterparts. Beyond mean estimation (e.g., Figure˜S11), they typically require constructing a grid over $\theta$. When the marginal density $\eta_{t}$ is not available in closed form (e.g., for the Student-$t$ prior), the Bayes-assisted version involves numerical integration. If computation is a concern, the Laplace prior offers a good compromise: it has heavier tails than the Gaussian while still admitting a closed-form expression for $\eta_{t}$.
Acknowledgments and Disclosure of Funding
Valentin Kilian is supported by the Clarendon Funds Scholarship. Stefano Cortinovis is supported by the EPSRC Centre for Doctoral Training in Modern Statistics and Statistical Machine Learning (EP/S023151/1). The authors thank the reviewers for their time and valuable feedback, especially the suggestion to incorporate a discussion on Type-I error control.
@@
The supplementary material is organised as follows. Appendix˜S1 gives additional background on strong laws and couplings, and confidence sequences. Appendix˜S2 states secondary results and their proofs. Appendix˜S3 presents the proofs of the main theorems and propositions. Appendix˜S4 extends the results to the multivariate setting. Appendix˜S5 gives specific expressions for the case of prediction-powered mean estimation. Appendix˜S6 details the experimental setup used in the main text. Appendix˜S7 presents further experiments. Finally, Appendix˜S8 discusses a parameter-free, non-assisted, AsympCS, and provides some additional comparisons.
For clarity, all sections, theorems, propositions, and lemmas in the supplementary material are prefixed with “S" to distinguish them from those in the main text.
Contents
@@
S1.1 Asymptotic theory of partial sums
S1.1.1 Iterated logarithm and Marcinkiewicz-Zygmund strong laws
@@
(Iterated Logarithm Law [1, Theorem 8.5.2]) Let $(Y_{t})_{t\geq 1}$ be i.i.d. random variables with zero mean and unit variance. Let $S_{t}=\sum_{i=1}^{t}Y_{i}$. Then,
$\underset{t\rightarrow\infty}{\lim\sup}\frac{|S_{t}|}{\sqrt{2t\log\log t}}=1\text{~~~a.s.},$
which implies
$\left|\frac{S_{t}}{t}\right|=O\left(\sqrt{\frac{\log\log t}{t}}\right)\text{~~~a.s. as }t\to\infty.$
Theorem S7.
(Marcinkiewicz-Zygmund strong law of large numbers [1, Theorem 2.5.12]) Let $(Y_{t})_{t\geq 1}$ be i.i.d. random variables with zero mean and $\mathbb{E}|Y_{1}|^{p}<\infty$ for some $1<p<2$. Let $S_{t}=\sum_{i=1}^{t}Y_{i}$. Then,
$\frac{S_{t}}{t^{1/p}}\rightarrow 0\text{~~~a.s. as }t\to\infty.$
S1.1.2 Strong approximations
The following strong invariance result, attributed to Komlós, Major and Tusnady (KMT) [2, 3] shows that the partial sums of i.i.d. random variables can be approximated almost surely by a Brownian motion path. The following theorem is from Csörgö and Hall [4, Theorem 3.2].
Theorem S8 (KMT strong coupling [2, 3]).
$S_{t}-B_{t}=o(t^{1/q})\text{~~~a.s.~as~}t\rightarrow\infty.$
KMT strong coupling has been extended by Einmahl [5] to random vectors (see also [6, Section B11]).
Theorem S9.
@@
$\Sigma^{-{1/2}}S_{t}-B_{t}=o(t^{1/q})\text{~~~a.s.~as~}t\rightarrow\infty.$
S1.2 Confidence sequences
S1.2.1 Confidence intervals vs. confidence sequences
Let $(X_{t})_{t\geq 1}$ be an observed data stream and let $\mu\in\mathbb{R}$ denote a fixed but unknown parameter (e.g., a mean). Write $\mathcal{F}_{t}=\sigma(X_{1:t})$ for the natural filtration, and let $\alpha\in(0,1)$ be a pre-specified error probability (so the confidence level is $1-\alpha$).
Fixed-time confidence intervals.
A (fixed-time) confidence interval (CI) for $\mu$ at time $t$ is an $\mathcal{F}_{t}$-measurable random set $\mathcal{C}_{\alpha,t}\subseteq\mathbb{R}$ such that
$\Pr\!\big(\mu\in\mathcal{C}_{\alpha,t}\big)\ \geq\ 1-\alpha.$
This guarantee is marginal in $t$: it holds for any chosen, deterministic $t$, but it need not be valid if $t$ is selected after looking at the data (e.g., by continual monitoring or a data-dependent stopping rule). In particular, for a general $\mathcal{F}_{t}$-stopping time $\tau$,
$\Pr\!\big(\mu\in\mathcal{C}_{\alpha,\tau}\big)\quad\text{can be }<\ 1-\alpha,$
unless the procedure is explicitly designed to be valid under optional stopping. Moreover, the family $(\mathcal{C}_{\alpha,t})_{t\geq 1}$ of fixed-time CIs need not be nested across $t$; disjoint intervals at different sample sizes can occur with positive probability (see Figure˜S4), illustrating the lack of any simultaneous-in-time guarantee.
Confidence sequences.
A confidence sequence (CS) at level $1-\alpha$ is a sequence of $\mathcal{F}_{t}$-measurable random sets $(\mathcal{C}_{\alpha,t})_{t\geq 1}$ such that
$\Pr\!\big(\mu\in\mathcal{C}_{\alpha,t}\ \text{for all }t\geq 1\big)\ \geq\ 1-\alpha.$
Equivalently,
$\Pr\!\Big(\sup_{t\geq 1}\mathbf{1}\{\mu\notin\mathcal{C}_{\alpha,t}\}=1\Big)\ \leq\ \alpha.$
The quantifier “for all tt” lies inside the probability, yielding uniform-in-time (a.k.a. anytime-valid) coverage. A key consequence is validity under arbitrary data-dependent stopping: for every (a.s. finite) stopping time $\tau$,
$\Pr\!\big(\mu\in\mathcal{C}_{\alpha,\tau}\big)\ \geq\ 1-\alpha.$
Thus CSs support continual monitoring and sequential decision-making without inflating error rates. Practically, CSs are typically wider than fixed-time CIs at the same $t$ (especially early on) because they control the maximum over all times; widths often shrink with $t$ and can approach classical rates up to iterated-logarithm factors.
Figure S4: Comparison of fixed-time confidence intervals (CIs) and a confidence sequence (CS) for data from $\mathcal{N}(0,1)$. Two fixed-time CIs at different sample sizes happen to be disjoint (highlighted), illustrating that marginal coverage at each $t$ does not imply simultaneous coverage over $t$. The CS is more conservative at small $t$, but its coverage holds uniformly over all $t$.
Early examples of CSs go back to sequential analysis [7, 8], and modern constructions often proceed via nonnegative supermartingales/test martingales and time-uniform concentration inequalities.
S1.2.2 Nonnegative supermartingale and Ville’s inequality
@@
$M=(M_{t})_{t\geq 1}$ is a nonnegative supermartingale (NSM) with respect to the filtration $(\mathcal{F}_{t})_{t\geq 1}$ if $M_{t}\geq 0$ a.s, $\mathbb{E}[M_{t}]<\infty$ for all $t\geq 1$, and
$\mathbb{E}[M_{t+1}\mid\mathcal{F}_{t}]\leq M_{t}\text{ a.s.}$
@@
Proposition S6 (Ville’s inequality [9]).
@@
Proposition S8 (Lemma B1, [10]).
Let $\{(M_{t}(\mu^{\prime}))_{t\in\mathbb{N}},\mu^{\prime}\in\mathbb{R}\}$ be a family of (super)martingales on a filtered probability space $(\Omega,\mathcal{A},(\mathcal{F}_{t})_{t\in\mathbb{N}},\Pr)$, indexed by $\mu^{\prime}$ in a measurable space $(\mathbb{R},\mathcal{B})$, such that
@@
    each $M_{t}(\mu^{\prime})$ is $\mathcal{F}_{t}\otimes\mathcal{B}$-measurable; and
@@
    each $\mathbb{E}\left[M_{t}(\mu^{\prime})\mid\mathcal{F}_{t-1}\right]$ is $\mathcal{F}_{t-1}\otimes\mathcal{B}$-measurable.
Let $\uppi$ be a finite measure on $(\mathbb{R},\mathcal{B})$ such that for all $n$,
$\Pr\otimes\uppi\text{-almost everywhere }M_{t}(\mu^{\prime})\geq 0,\quad\text{ or }\quad\mathbb{E}_{\mu^{\prime}\sim\uppi}\mathbb{E}\left[\left|M_{t}(\mu^{\prime})\right|\right]<\infty$
Then the mixture $\left(\tilde{M}_{t}\right)_{t\in\mathbb{N}}$, where $\tilde{M}_{t}=\mathbb{E}_{\mu^{\prime}\sim\uppi}M_{t}(\mu^{\prime})$, is also a (super)martingale.
This is useful as it leads to the method of mixtures : if we have a family of nonnegative supermartingale (say) of the form $M_{t}(\mu^{\prime})$ for $\mu^{\prime}\in\mathbb{R}$ which satisfy conditions 1 and 2 above and a mixture distribution $\uppi$ satisfying the assumptions of Proposition˜S8, then we can conclude that $\int_{\mu^{\prime}\in\mathbb{R}}M_{t}(\mu^{\prime})d\uppi(\mu^{\prime})$ is also a supermartingale, and thus Ville’s inequality gives for any $\alpha\in(0,1)$
$\Pr\left(\int_{\mu^{\prime}\in\mathbb{R}}M_{t}(\mu^{\prime})d\uppi(\mu^{\prime})\leq\frac{1}{\alpha}\text{ for all }t\geq 1\right)\geq 1-\alpha.$ (S28)
The method of mixtures dates back at least to Ville [9] and was developed in the context of sequential analysis by Wald [11]. It was then systematised and popularised by Darling and Robbins in the late 1960s, by Robbins and Siegmund in a series of papers culminating in [7], and by Lai [8]. The method of mixtures has found many applications, including confidence sequences [8, 12, 13, 14, 15], PAC-Bayes analysis [16, 10], anytime-valid testing [17], and A/B testing [18], to name but a few.
@@
The following proposition follows from KMT strong approximation (see Theorem˜S8). It will be used in the proofs of Theorem˜3 and Proposition˜2.